# PAE MobiLLM: Privacy-Aware and Efficient LLM Fine-Tuning on the Mobile Device via Additive Side-Tuning

### v1

@@ Abstract.
There is a huge gap between numerous intriguing applications fostered by on-device large language model (LLM) fine-tuning (FT) from fresh mobile data and the limited resources of a mobile device. While existing server-assisted methods (e.g., split learning or side-tuning) may enable LLM FT on the local mobile device, they suffer from heavy communication burdens of activation transmissions, and may disclose data, labels or fine-tuned models to the server. To address those issues, we develop PAE MobiLLM, a privacy-aware and efficient LLM FT method which can be deployed on the mobile device via server-assisted additive side-tuning. The proposed PAE MobiLLM is featured by (i) notable communication and computing efficiency improvement; (ii) keeping data, true labels, and fine-tuned full models on the local mobile device; (iii) fast side-tuning on the server side. Briefly, PAE MobiLLM divides LLM fine-tune workloads between the mobile device and the server, i.e., the mobile device merely retains a frozen backbone model, while the computation-intensive trainable side-network is offloaded to the server. To further accelerate FT convergence and improve computing efficiency, PAE MobiLLM integrates activation caching on the server side, which allows the server to reuse historical activations and saves the mobile device from repeatedly computing forward passes for the recurring data samples. Besides, to reduce communication cost, PAE MobiLLM develops a one-token (i.e., “pivot” token) activation shortcut that transmits only a single activation dimension instead of full activation matrices to guide the side network tuning. Last but not least, PAE MobiLLM introduces the additive adapter side-network design which makes the server train the adapter modules based on device-defined prediction differences rather than raw ground-truth labels. In this way, the server can only assist device-defined side-network computing, and learn nothing about data, labels or fine-tuned models. Extensive experimental results demonstrate PAE MobiLLM’s superiority.
LLM fine-tuning, Mobile device, Side-tuning, Computing and communication efficiency, Privacy awareness.
Both authors contributed equally to this research.
ccs: Computing methodologies Distributed artificial intelligence
1. Introduction
@@
Large language models (LLMs) such as GPT (2), RoBERTa (17) and LLaMA (27) have recently demonstrated remarkable capabilities across various NLP tasks, including language representation (22), cross-task reasoning (34), and context-aware generation (18). This success stems from their two-stage adaptation paradigm: pretraining on large-scale public corpora followed by domain-specific fine-tuning (FT). While effective in centralized cloud environments, this conventional approach faces fundamental challenges in mobile computing scenarios, where user data is continuously generated across distributed mobile devices (e.g., smartphones and wearables), highly privacy-sensitive (e.g., health monitoring records and personal chat histories, and subject to stringent latency demands for intelligent services (15; 33). By adapting a local LLM with fresh data on the mobile device, on-device fine-tuning is emerging as both a practical necessity and a strategic shift toward personalized intelligence without compromising user privacy.
@@
Despite those promising applications, on-device LLM FT faces significant computing resource challenges. Because LLMs scale to billions of parameters, both the computational and memory requirements for fine-tuning an LLM can be prohibitively high for existing mobile devices’ hardware. Beyond storing the full model parameters, the predominant LLM training algorithm necessitates substantial memory for intermediate activations, gradients, and optimizer states. For example, fine-tuning a 7B-parameter LLM typically requires over 70 GB of GPU memory, far exceeding the capacity of most mobile devices, which often have 4–12 GB of RAM (32). A single training epoch for such an LLM could require hours of processing on mobile GPUs (3), which means full on-device fine-tuning cycles will take multiple days. This memory usage and latency fundamentally conflict with practical deployment scenarios requiring continuous model adaptation.
@@
To enable LLM FT on a computing resource constrained mobile device, researchers are exploring to get some help from the high-performance computing server and collaborate the mobile device and the server via wireless communications supported by current mobile edge computing infrastructures. An intuitive solution is to integrate split learning (11) with PEFT, i.e., training shallow layers on the mobile device while offloading deeper layers to the server as shown in Fig. 1(a). Yet, the sequential processing and bidirectional communications in this approach require fast wireless access and strict synchronization. Another device-server collaboration LLM FT approach is MobiLLM (16), which introduces server-assisted parallel side-tuning. Instead of splitting the model vertically by layers, MobiLLM decouples the trainable modules (hosted on the server) from the frozen backbone model (retained on the mobile device) as shown in Fig. 1(b). During LLM FT, the mobile device executes forward passes on the backbone model and transmits intermediate activations to the server. The server computes gradients and updates the side network via backpropagation, which helps the mobile device bypass the computing and memory intensive gradient calculations. By eliminating backpropagation, MobiLLM enables the LLM FT of billion-parameter models (e.g., OPT-1.3B) on the mobile device within 4.5 GB memory.
@@
Although the pioneering device-server collaboration framework relieves the memory concerns of mobile devices, it incurs new challenges for the practical deployment of LLM FT on the mobile device. (i) Iterative on-device computation cost during LLM FT: Existing device-server collaboration designs require the mobile devices to iteratively compute the activations across all training epochs. That induces enormous cumulative computation costs on the resource-constrained mobile device, which grows proportionally with LLM FT duration. (ii) Extra communication burden of activation transmissions: Both split learning and MobiLLM have to transfer full-dimension activations/gradients iteratively to the server, which incurs extra communication burdens. Especially when processing large batches with long sequence inputs, the communication cost may be too high to execute LLM FT with low-speed access or time-varying channels in wireless edge networks. (iii) Label or fine-tuned model exposure to the server: One major objective of on-device LLM FT is to keep data, labels, and fine-tuned models on the local mobile device. However, at the cost of server’s computing assistance, current device-server collaboration methods have the risk to disclose the true labels or fine-tuned models to the server. For instance, MobiLLM necessitates transmitting supervisory labels from the device to the server for parallel side-network training. Besides, the server can easily infer the fine-tuned model by adding up the tunable side network known by the server and the public frozen backbone model.
To address those issues above under device-server collaboration framework, in this paper, we propose PAE MobiLLM, a privacy-aware and efficient LLM FT approach, which can be deployed on the mobile device with the server’s help while maintaining data, labels and fine-tuned models on the local device [^1] Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels, and fine-tuned models on the local device, which is similar to the data privacy protection definition in Google’s federated learning (FL) work  (19). Besides, the same as FL, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, such as differential privacy (DP) mechanisms  (10) (e.g., DP noises added on gradients, input, output or objective functions), secure multiparty computing (9), etc.. Our salient contributions are summarized as follows:
  • We develop an activation caching mechanism that allows the server to store intermediate activations from previously processed data samples. This not only accelerates the side-network training on the server side, but also enables the device to skip redundant forward propagation on the frozen backbone for recurring samples, reducing both computational and transmission costs.
  • We observe that a single activation dimension rather than the full activation matrix is good enough to guide adapter side-tuning of transformer-based models. Thus, we introduce a one-token activation transmission mechanism for shortcut connections between the mobile device and the server, which helps to significantly reduce communication costs and latency.
  • We design an additive adapter side-network, which can leverage the server’s computing capability to facilitate fine-tuning without disclosing ground-truth labels. The proposed additive adapter side-network consists of cascaded lightweight adapter modules that fuse model predictions via output difference-guided interpolation, which thereby eliminates the need for direct label transmission from the mobile device to the server.
  • We implement PAE MobiLLM on various mobile devices and development kits, including NVIDIA Jetson Xavier and Huawei MateBook, and evaluate its performance on sub-billion and billion-scale LLMs across diverse tasks and system settings. Experimental results demonstrate that compared with the second best LLM FT method on the mobile device, PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × reduction in convergence time, while keeping data, true labels and fine-tuned models on the mobile device.
The remainder of the paper is organized as follows. Section II gives an introduction to LLM fine-tuning. Section III analyzes the limitations of existing works and presents our motivation. Section IV elaborates on our PAE MobiLLM design. Section V specifies experimental configurations. Section VI presents the performance evaluation. Section VII discusses our future work and Section VIII finally concludes the paper.
2. Preliminaries
2.1. Parameter- and Memory-Efficient FT
@@
The memory footprint of LLM fine-tuning arises from three core components: model parameters, intermediate activations, and optimizer states. Extensive research has focused on optimizing these aspects, with Parameter-efficient Fine-tuning (PEFT) methods emerging as a prominent approach to reduce optimizer state memory and gradient computation costs by selectively updating sparse subsets of parameters while maintaining performance comparable to full fine-tuning. Techniques such as Adapter (12), which inserts bottleneck modules between transformer layers, LoRA (13), which decomposes weight updates using low-rank matrices, and BitFit (35), which exclusively trains bias terms, exemplify this paradigm. Further integration of model compression with PEFT—for instance, QLoRA (6)’s application of 4-bit quantization to the backbone network—has enabled significant reductions in parameter memory footprint. However, these approaches remain constrained by activation memory overhead as their trainable parameters remain embedded within the original model architecture, necessitating full backward passes through all layers. To address this limitation, recent innovations like ladder side-tuning (LST) (26) and LoSA (20) propose decoupling trainable components into a lightweight side network, creating a backpropagation highway that bypasses activation storage for the original model. Despite such advancements, practical deployment on resource-constrained edge devices faces unresolved challenges: low-bit quantization often demands specialized hardware support, and even memory-efficient designs like LST still exceed typical edge device capacities (16). These facts limit their feasibility in extremely resource-constrained environments.
Figure 1. Device-server collaborative LLM FT.
2.2. Device-Server Collaborative LLM FT
@@
To enable device-side LLM adaptation while preserving private data, researchers are exploring on-device fine-tuning techniques. Given mobile hardware constraints, a common strategy involves partitioning the LLM and offloading part of the computation/storage burdens to adjacent servers or devices. Particularly, the U-shaped split learning (23) structure retains and updates only the few input and output transformer layers on the mobile device while delegating intermediate layer training to servers, ensuring that user data and labels remain local. As shown in Fig.  1(a), the training framework integrates U-shaped split learning with LoRA in a device-server collaborative setup, where activations of split layers are transmitted during forward propagation, and corresponding gradients are exchanged during backpropagation. When extended to multi-device scenarios, the LLM is further divided into layer-wise sub-models across devices, with forward/backward computations executed sequentially in a relay fashion (5). However, this layer-wise partitioning enforces strict sequential execution, and requires each device to update partial trainable parameters. In contrast, MobiLLM (16) introduces an innovative fine-tuning framework that decouples LLM fine-tuning into a frozen backbone model and a trainable side network, where the server hosts the adapter-based side network parallel to the device-side backbone (Fig. 1(b)) This framework maintains memory-efficient forward passes and preserves the original LLM functionality on resource-constrained devices, while offloading computation-intensive backpropagation to high-performance servers. Crucially, this parallel device-server training allows the device to execute lightweight inference through the frozen backbone while the server simultaneously computes gradients via the side network, which achieves efficient collaborative training without exposing raw input data.
2.3. “Pivot” Token in LLM architectures
Modern pre-trained LLMs predominantly adopt the Transformer architecture, where self-attention mechanisms capture dependencies across all tokens in an input sequence, enabling robust sentence-level comprehension. Fig. LABEL:fig:token_flow illustrates the activation dimension flow in a binary sentiment analysis task with a 3-token input sequence. After passing through the embedding layer, the input sequence is mapped to 3 tokens. The first transformer layer projects them to a shape of $[3,H]$, which remains unchanged in all subsequent transformer layers, where $H$ is the hidden size. At the final layer, a classification head projects the activations to $[3,2]$, requiring aggregation into a $[1,2]$ decision vector. Different architectures employ distinct token selector $S(\cdot)$ to extract global sentence representations, reflecting their pretraining objectives: For Auto-regressive models (e.g., GPT (2)) trained via left-to-right language modeling, they implicitly encode global context into the last token’s activation. The final token’s hidden state naturally accumulates bidirectional dependencies through stacked self-attention layers, serving as the de facto sentence representation. For auto-encoding models (e.g., BERT (7)) utilizing bidirectional context modeling, they prepend a dedicated [CLS] token to the input sequence. During pretraining, the [CLS] token’s embeddings are explicitly optimized to aggregate global sentence features through next-sentence prediction objectives, making it the canonical choice for downstream classification tasks. In conclusion, most LLMs designate a pivotal token ([CLS] in BERT, final token in GPT) to encode global context, making this token critical for the fine-tuning process compared to other positional dimensions.
Figure 2. Activation dimension changes in LLMs.
3. Motivation
@@
Observation 1: Iterative computation on the mobile device. LLM FT tasks typically demand dozens to hundreds of training epochs for convergence, requiring iterative training on the same set of local samples. While device-server collaborative frameworks offload partial computations to the server, mobile devices remain burdened by iterative local computation. For instance, in the split learning framework, the mobile device must recompute and transmit the intermidiate activations in every epoch—even for identical input samples—due to continuous parameter updates on the device-side model. MobiLLM eliminates device-side backpropagation by maintaining a frozen pretrained backbone, but it still enforces re-computation and transmission of static intermediate activations across training epochs by default. Table 1 compares communication and computational costs of device-only BitFit, split-learning BitFit, and MobiLLM. Since the required number of epochs in a fine-tuning session usually depends on dynamic factors like dataset size and hyperparameter configurations, prolonging on-device participation could make resource planning on memory/energy-constrained devices inherently unstable and inefficient.
Table 1. Communication and computational costs (On Xavier, task: MRPC@OPT-1.3B, batch size=8, seq len=256).
Method
Per-Batch
Comm. (MB)
Per-Epoch
Comp. (PFLOPS)
Total Comp.
BitFit-local - 18.4 208.3
BitFit-SL 33.4 1.5 17.4
MobilLM 200.4 9.5 124.5
@@
Observation 2: High device-server communication burden. Current collaborative LLM FT methods require frequent device-server communications for exchanging intermediate activations or gradients—bidirectional in split learning (activations and gradients) and unidirectional in MobiLLM (activations only). This results in substantial transmission loads, especially for large models processing long sequences or large batches. For instance, in U-shaped split learning, fine-tuning OPT-1.3B with a batch size of 64 and token length of 512 requires transmitting up to 0.52 GB of data for just one batch. Despite advanced future wireless accessing technologies (e.g., 6G), communications remain a critical issue given the exponential growth of LLM parameters and edge deployment demands. However, a key architectural insight is often overlooked: each transformer layer inherently generates a “pivot” token that aggregates global semantic information and guides the model’s final decision. Existing methods transmit all token-level activations $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size) by default, despite the pivot token’s compact representation $[B,H]$ being sufficient for LLM adaptation. That may offers a path to orders-of-magnitude efficiency gains - an opportunity yet unexplored in collaborative LLM FT paradigms.
Figure 3. An overview of the PAE MobiLLM system.
@@
Observation 3: Unaware of label and fine-tuned model privacy. Existing device-server collaborative FT frameworks allow local retention of raw input data while requiring transmission of ground-truth labels to the server for loss calculation and partial parameter updates, e.g., updating deep layers in split learning or side-network parameters in MobiLLM. It may introduce systemic privacy vulnerabilities for label-sensitive applications, such as generative content creation or medical diagnostics. For instance, healthcare labels often contain highly sensitive information (e.g., disease diagnoses or genetic conditions). Moreover, these designs may inadvertently expose fine-tuned models. In MobiLLM, the server can combine the tuned side-network with the publicly known backbone model to replicate the device-specific fine-tuned LLM. Similarly, in split learning, although servers only update partial tunable parameters, they could maintain synchronized copies of client-side shallow layer initializations and mirror parameter updates to assemble the full model. While the U-shaped split learning variants (shown in Fig. 1(b)) keep labels locally, they fundamentally double the bidirectional activation/gradient transmission cycles, incurring excessive communication cost and increasing training latency as discussed in Observation 2.
4. PAE MobiLLM Design
4.1. PAE MobiLLM Overview
@@
To address the observed issues, we propose PAE MobiLLM, a novel privacy-aware and efficient LLM FT design for the mobile device. PAE MobiLLM follows the device-server collaboration framework but effectively improves both computational and communication efficiency while ensuring data, label, i.e., $(\text{Input}_{x},\text{Label}_{y})$, and the fine-tuned model never leave the device. Briefly, as shown in Fig. 3, PAE MobiLLM assigns a frozen pre-trained model on the mobile device and lets it perform forward propagation only during the first epoch when batching over the local dataset. During this process, the mobile device forwards two types of information to the server: (1) the activation values ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) between transformer layers, which will be compressed by a token selector $S(\cdot)$ to reduce communication cost before transmissions; (2) the adjustment term $\Delta y$, as a linear calculation result of the backbone network’s forward inference result $y_{\text{pre}}$, the ground-truth label $\text{Label}_{y}$, and a locally generated random nonce $R$. PAE MobiLLM allocates all trainable modules to the server. During the first epoch, the server updates model parameters using ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) received from the device, while caching these ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) to construct a side-tuning guidance database. After the first epoch, the mobile device is freed, and the server will train the side-network iteratively until its convergence based on the cached activations. We will detail the key components, their functionalities and the procedure of PAE MobiLLM in the subsections below.
4.2. Key Computing-Efficient Components
4.2.1. Adapter Side-Tuning Network
@@
Aiming to improve computing efficiency, PAE MobiLLM follows the device-server collaboration framework to offload the computing-intensive workload to the server. As shown in Fig.  3, PAE MobiLLM lets the mobile device only retain the frozen pre-trained model for forward inference and offloads the training and updating of the side-network to the server. In the side-tuning network, the trainable modules refine transformer-layer outputs from the backbone network, and enhance the representation quality for LLM FT. In PAE MobiLLM, for a pre-trained model with $L$ transformer layers, we construct $L$ adapter modules, where the $i$-th adapter module corresponds to the $i$-th transformer layer. The input and output dimensions of each adapter module match the hidden size of the transformer layers, ensuring seamless optimization of the backbone’s representations. The fine-tuning operations of the $i$-th layer is outlined in Fig.  4(a).
@@
(2) $h_{S_{i}}=S_{i\_in}+\sigma(S_{i\_in}\cdot W_{\text{down}})W_{\text{up}}.$
@@
4.2.2. Activation Caching Mechanism on the Server
@@
LLM FT typically requires multiple iterations of training, while iterative on-device training for activation updates will raise huge computational cost for the mobile device for either split learning or MobiLLM as discussed in Sec. 3. However, under the proposed PAE MobiLLM design, we observe that the backbone network and the side network work independently, and the backbone network on the mobile device remains frozen as shown in Fig. 3. This means that for the same data samples, activations remain unchanged across iterations.
@@
Inspired by this observation, we propose an activation caching mechanism on the server. As shown in the rightmost part of Fig. 3, during the first epoch, the mobile device transmits intermediate activations ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) and the training target $\Delta\mathbf{y}$ to the server, while the server trains the side network and caches the received activations and $\Delta\mathbf{y}$. After the first epoch, since Inputx is the same and the backbone network is frozen on the mobile device, activations stay unchanged. So, the mobile device is freed. The server will train the side network iteratively until its convergence based on the cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$). The activation caching mechanism accelerates the side-tuning since the activation values from the backbone network are the same every iteration and the server can continuously conduct side-tuning using cached values without waiting for any activation updates.
Besides, due to activation caching, the mobile device only needs to compute the forward pass once and transmit the activation values to the server in the first epoch. That effectively improves computing efficiency and significantly reduces communication cost for the mobile device.
4.3. Key Communication-Efficient Components

### v2

@@ Abstract
There is a huge gap between numerous intriguing applications fostered by on-device large language model (LLM) fine-tuning (FT) from fresh mobile data and the limited resources of a mobile device. While existing server-assisted methods (e.g., split learning or side-tuning) may enable LLM FT on the local mobile device, they suffer from heavy communication burdens of activation transmissions, and may disclose data and labels to the server. To address those issues, we develop PAE MobiLLM, a privacy-aware and efficient LLM FT method which can be deployed on the mobile device via server-assisted additive side-tuning. To further accelerate FT convergence and improve computing efficiency, PAE MobiLLM integrates activation caching on the server side, which allows the server to reuse historical activations and saves the mobile device from repeatedly computing forward passes for the recurring data samples. Besides, to reduce communication cost, PAE MobiLLM develops an activation shortcut that transmits only the token involved in the loss calculation instead of full activation matrices to guide the side network tuning. Last but not least, PAE MobiLLM introduces the additive adapter side-network design which makes the server train the adapter modules based on device-defined prediction differences rather than raw ground-truth labels. In this way, the server can only assist device-defined side-network computing, and learn nothing about data and labels. Extensive experimental results demonstrate PAE MobiLLM’s superiority.
Introduction
@@
Large language models (LLMs) like GPT (Brown et al. 2020) and LLaMA (Touvron et al. 2023) have recently demonstrated remarkable capabilities across various NLP tasks (Ren et al. 2024; Ye 2024; Brown et al. 2020) via pretraining on public corpora followed by domain-specific fine-tuning (FT). Although effective in centralized cloud environments, this paradigm faces fundamental challenges in mobile computing scenarios, where user data are continuously generated across distributed mobile devices (e.g., smartphones and wearables), highly privacy-sensitive (e.g., health monitoring records and personal chat histories), and subject to stringent latency demands for intelligent services (Li et al. 2021; Ye et al. 2020). How to enable LLM FT on the mobile device without compromising data privacy becomes an essential and popular research topic.
@@
Despite those promising applications, on-device LLM FT faces significant computing resource challenges. Because LLMs scale to billions of parameters, both the computational and memory requirements for fine-tuning an LLM can be prohibitively high for existing mobile devices’ hardware. Beyond storing the full model parameters, the predominant LLM training algorithm necessitates substantial memory for intermediate activations, gradients, and optimizer states. For example, fine-tuning a 7B-parameter LLM typically requires over 70 GB of GPU memory, far exceeding the capacity of most mobile devices, which often have 4–12 GB of RAM (Xu et al. 2024). A single training epoch for such an LLM could require hours of processing on mobile GPUs (Cai et al. 2023), which means full on-device fine-tuning cycles will take multiple days. This memory usage and latency fundamentally conflict with practical deployment scenarios requiring continuous model adaptation.
@@
To enable LLM FT on a computing resource constrained mobile device, researchers are exploring to get some help from the high-performance computing server and collaborate the mobile device and the server via wireless communications supported by current mobile edge computing infrastructures. An intuitive solution is to integrate split learning (Gupta and Raskar 2018) with PEFT, i.e., training shallow layers on the mobile device while offloading deeper layers to the server. Yet, the sequential processing and bidirectional communications in this approach require fast wireless access and strict synchronization. Another device-server collaboration LLM FT approach is MobiLLM (Li et al. 2025), which introduces server-assisted parallel side-tuning. Instead of splitting the model vertically by layers, MobiLLM decouples the trainable modules (hosted on the server) from the frozen backbone model (retained on the mobile device). During LLM FT, the mobile device executes forward propagation on the backbone model and transmits intermediate activations to the server. The server computes gradients and updates the side network via backpropagation, which helps the mobile device bypass the computing and memory intensive gradient calculations. By eliminating backpropagation, MobiLLM enables the LLM FT of billion-parameter models (e.g., OPT-1.3B) on the mobile device within 4.5 GB memory.
@@
Although the pioneering device-server collaboration framework relieves the memory concerns of mobile devices, it incurs new challenges for the practical deployment of LLM FT on the mobile device. (i) Iterative on-device computation cost during LLM FT: Existing device-server collaboration designs require the mobile devices to iteratively compute the activations across all training epochs. That induces enormous cumulative computation costs on the resource-constrained mobile device, which grows proportionally with LLM FT duration. (ii) Extra communication burden of activation transmissions: Both split learning and MobiLLM have to transfer full-dimension activations/gradients iteratively to the server, which incurs extra communication burdens. Especially when processing large batches with long sequence inputs, the communication cost may be too high to execute LLM FT with low-speed access or time-varying channels in wireless edge networks. (iii) Raw Label exposure to the server: One major objective of on-device LLM FT is to keep data and labels on the local mobile device. However, at the cost of server’s computing assistance, current device-server collaboration methods have the risk to disclose the true labels to the server. For instance, MobiLLM necessitates transmitting supervisory labels from the device to the server for side-network training.
To tackle those issues, in this paper, we propose PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via device-server collaboration. PAE MobiLLM keeps data and labels local, while delegating memory/compute-intensive operations to an edge server via additive side-tuning. Our contributions are summarized as follows: (i) We design an additive adapter side-network, which can leverage the server’s computing capability to facilitate fine-tuning without disclosing ground-truth labels. The proposed additive adapter side-network consists of cascaded lightweight adapter modules that fuse model predictions via output difference-guided interpolation, which thereby eliminates the need for direct label transmission from the mobile device to the server. (ii) We introduce a token selection method that transmits only critical backbone-derived activations to supervise the side-network training at the server, thereby significantly reducing the communication workload. (iii) We develop an activation caching mechanism that allows the server to store intermediate activations from previously processed data samples. This not only accelerates the side-network training on the server side, but also enables the device to skip redundant forward propagation on the frozen backbone for recurring samples, reducing both computational and transmission costs. (iv) We implement PAE MobiLLM on various mobile devices and evaluate its performance on sub-billion and billion-scale LLMs across diverse tasks and system settings. Experimental results demonstrate that compared to the second best LLM FT method on the mobile device, PAE MobiLLM achieves at least a $1.79\times$ reduction in device-side computation, a $17.06\times$ reduction in communication cost and a $5.25\times$ reduction in convergence time, while keeping data and true labels private.
Preliminaries
Parameter- and Memory-Efficient FT
@@
The memory footprint of LLM fine-tuning arises from three core components: model parameters, intermediate activations, and optimizer states. Extensive research has focused on optimizing these aspects, with Parameter-efficient Fine-tuning (PEFT) methods emerging as a prominent approach to reduce optimizer state memory and gradient computation costs by selectively updating sparse subsets of parameters while maintaining performance comparable to full fine-tuning. Techniques such as Adapter (Houlsby et al. 2019), which inserts bottleneck modules between transformer layers, LoRA (Hu et al. 2022), which decomposes weight updates using low-rank matrices, and BitFit (Zaken, Ravfogel, and Goldberg 2021), which exclusively trains bias terms, exemplify this paradigm. However, these approaches remain constrained by activation memory overhead as their trainable parameters remain embedded within the original model architecture, necessitating full backward passes through all layers. To address this limitation, recent innovations like ladder side-tuning (LST) (Sung, Cho, and Bansal 2022) and LoSA (Mercea et al. 2024) propose decoupling trainable components into a lightweight side network, creating a backpropagation highway that bypasses activation storage for the original model.
Device-Server Collaborative LLM FT
@@
To enable device-side LLM adaptation while preserving private data, researchers are exploring on-device fine-tuning techniques. Given mobile hardware constraints, a common strategy involves partitioning the LLM and offloading part of the computation/storage burdens to adjacent servers or devices. Particularly, the U-shaped split learning (Ronneberger, Fischer, and Brox 2015) structure retains and updates only the few input and output transformer layers on the mobile device while delegating intermediate layer training to servers, ensuring that user data and labels remain local. When extended to multi-device scenarios, the LLM is further divided into layer-wise sub-models across devices, with forward/backward computations executed sequentially in a relay fashion (Chen et al. 2023). However, this layer-wise partitioning enforces strict sequential execution, and requires each device to update partial trainable parameters. In contrast, MobiLLM (Li et al. 2025) introduces an innovative framework that decouples LLM FT into a frozen backbone model and a trainable side network, where the server hosts the adapter-based side network parallel to the device-side backbone. This framework maintains memory-efficient forward passes and preserves the original LLM functionality on resource-constrained devices, while offloading computation-intensive backpropagation to high-performance servers. Crucially, this parallel device-server training allows the device to execute lightweight inference through the frozen backbone while the server simultaneously computes gradients via the side network, which achieves LLM FT without exposing raw input data.
Figure 1: An overview of the PAE MobiLLM system.
Motivation
@@
Observation 1: Iterative computation on the mobile device. LLM FT tasks typically demand dozens to hundreds of training epochs for convergence, requiring iterative training on the same set of local samples. While device-server collaborative frameworks offload partial computations to the server, mobile devices remain burdened by iterative local computation. For instance, in the split learning framework, the mobile device must recompute and transmit the intermidiate activations in every epoch—even for identical input samples—due to continuous parameter updates on the device-side model. MobiLLM eliminates device-side backpropagation by maintaining a frozen pretrained backbone, but it still enforces re-computation and transmission of static intermediate activations across training epochs by default.
@@
Observation 2: High device-server communication burden. Current collaborative LLM FT methods require frequent device-server communications for exchanging intermediate activations or gradients—bidirectional in split learning (activations and gradients) and unidirectional in MobiLLM (activations only). This results in substantial transmission loads, especially for large models processing long sequences or large batches. For instance, in U-shaped split learning, fine-tuning OPT-1.3B with a batch size of 64 and token length of 512 requires transmitting up to 0.52 GB of data for just one batch. Despite advanced future wireless accessing technologies (e.g., 6G), communications remain a critical issue given the exponential growth of LLM parameters and edge deployment demands.
@@
Observation 3: Unaware of label privacy. Existing device-server collaborative FT frameworks allow local retention of raw input data while requiring transmission of ground-truth labels to the server for loss calculation and partial parameter updates, e.g., updating deep layers in split learning or side-network parameters in MobiLLM. It may introduce systemic privacy vulnerabilities for label-sensitive applications, such as generative content creation or medical diagnostics. For instance, healthcare labels often contain highly sensitive information (e.g., disease diagnoses or genetic conditions), and in generation tasks, the labels can even be part of the raw input data. While the U-shaped split learning variants keep labels locally, they fundamentally double the bidirectional activation/gradient transmission cycles, incurring excessive communication cost and increasing training latency as discussed in Observation 2.
PAE MobiLLM Design
PAE MobiLLM Overview
@@
To address the observed issues, we propose PAE MobiLLM, a novel privacy-aware and efficient LLM FT design for the mobile device. PAE MobiLLM follows the device-server collaboration framework but effectively improves both computational and communication efficiency while ensuring data, label, i.e., $(\text{Input}_{x},\text{Label}_{y})$ never leave the device. Briefly, as shown in Fig. 1, PAE MobiLLM assigns a frozen pre-trained model on the mobile device and lets it perform forward propagation only during the first epoch when batching over the local dataset. During this process, the mobile device forwards two types of information to the server: (1) the activation values ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) between transformer layers, which will be compressed by a token selector $S(\cdot)$ to reduce communication cost before transmissions; (2) the adjustment term $\Delta y$, as a linear calculation result of the backbone network’s forward inference result $y_{\text{pre}}$, the ground-truth label $\text{Label}_{y}$. PAE MobiLLM allocates all trainable modules to the server. During the first epoch, the server updates model parameters using ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) received from the device, while caching these ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) to construct a side-tuning guidance database. After the first epoch, the mobile device is freed, and the server will train the side-network iteratively until its convergence based on the cached activations. We will detail the key components, their functionalities and the procedure of PAE MobiLLM in the subsections below.
Key Computing-Efficient Components
Adapter Side-Tuning Network
@@
Aiming to improve computing efficiency, PAE MobiLLM follows the device-server collaboration framework to offload the computing-intensive workload to the server. As shown in Fig.  1, PAE MobiLLM lets the mobile device only retain the frozen pre-trained model for forward inference and offloads the training and updating of the side-network to the server. In PAE MobiLLM, for a pre-trained model with $L$ transformer layers, we construct $L$ adapter modules, where the $i$-th adapter module corresponds to the $i$-th transformer layer. The input and output dimensions of each adapter module match the hidden size of the transformer layers, ensuring seamless optimization of the backbone’s representations. The fine-tuning operations of the $i$-th layer is outlined in Fig.  2.
@@
$h_{S_{i}}=S_{i\_in}+\sigma(S_{i\_in}\cdot W_{\text{down}})W_{\text{up}}.$ (2)
@@
Activation Caching Mechanism on the Server
@@
Under the proposed PAE MobiLLM design, we observe that the backbone network and the side network work independently, and the backbone network on the mobile device remains frozen as shown in Fig. 1. This means that for the same data samples, activations remain unchanged across iterations.
@@
Inspired by this observation, we propose an activation caching mechanism on the server. As shown in the rightmost part of Fig. 1, during the first epoch, the mobile device transmits intermediate activations ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) and the training target $\Delta\mathbf{y}$ to the server, while the server trains the side network and caches the received activations and $\Delta\mathbf{y}$. After the first epoch, since Inputx is the same and the backbone network is frozen on the mobile device, activations stay unchanged. So, the mobile device is freed. The server will train the side network iteratively until its convergence based on the cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$). Due to the activation caching mechanism, the mobile device only needs to compute the forward pass once and transmit the activation values to the server in the first epoch. That effectively improves computing efficiency and significantly reduces communication cost.
Key Communication-Efficient Components